OpenStudy (anonymous):

x over 12 = 6 over 7

OpenStudy (saifoo.khan):

\[\frac{x}{12} = \frac67 \]Cross multiply. 7x = 6 * 12 solve for x.

OpenStudy (anonymous):

x=72/7